Circular No. 7846

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bn IN UGC 6527 M. Armstrong, Rolvenden, England, reports his discovery of an ap-

parent supernova (mag 17.0) on an unfiltered CCD image (limiting mag 18.5) taken on Mar. 8.180 UT in the course of the U.K. Nova/Supernova patrol, and confirmed at mag 16.3 in an earlier image taken on Feb. 15.948. The new object is located at α = 11^{h}32^{m}39^{s}.04, δ = +52^{o}56′54′′.2 (equinox 2000.0), or ≈ 10′′ west and 5′′ south of the center of UGC 6527. SN 2002bn is not present on Armstrong’s images from 2000 Dec. 30 (limiting mag 19.0) or 2001 Dec. 10 and 22, or on Palomar Sky Survey images taken on 1993 Apr. 25 (limiting red mag 20.8) and 1994 Apr. 19 (limiting blue mag 22.5).

SUPERNOVAE 2002bf, 2002bg, 2002bi, AND 2002bk T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report: “A spectrum (range 370–750 nm) of SN 2002bf (cf. IAUC 7836), obtained by P. Berlind on Mar. 6.21 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at or before maximum. Adopting the NED recession velocity of 7254 km/s for the host galaxy, the supernova expansion velocity is ∼ 15 200 km/s for Si II (rest 635.5 nm) and 19 400 for Ca II (rest 395.1 nm), implying an age of ≈ two weeks before maximum, based upon comparison with values presented by Jha et al. (1999, Ap.J. Supp. 125, 73). However, the spectral-feature age (cf. IAUC 7774) of the supernova is 0 ± 2 days before maximum light. Given the epoch of discovery (Feb. 22.3), and the general appearance of the spectrum, the latter value is probably more accurate. In that case, the high expansion velocities are unusual.”

A. V. Filippenko and R. Chornock, University of California, Berkeley; D. C. Leonard, University of Massachusetts, Amherst; and E. C. Moran, Wesleyan University, write: “Inspection of CCD spectra (range 390–880 nm) obtained on Mar. 7 UT with the Keck-I telescope reveals that SN 2002bf is of type Ia, within a few days of maximum brightness; the expansion velocity of the ejecta, measured from the minimum in the Si II 635.5 nm trough, is ∼ 15 000 km/s — a value that is unusually high for a type-Ia supernova at this phase. SN 2002bg (IAUC 7836) is also of type Ia, 14 ± 2 days past maximum brightness, in agreement with the conclusions of Matheson et al. (IAUC 7844). SN 2002bi (IAUC 7837) is, again, clearly of type Ia, roughly 1 month past maximum. Similarly, SN 2002bk (IAUC 7842) is of type Ia, ∼ 5 weeks past maximum; the host galaxy recession speed is 17 000 km/s, estimated from the supernova spectrum itself.”

2002 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green